A boy with a mass 50.0 kg walks up the stairs from the ground to the top of a spiral water slide in 1.00 minute. The height of the slide is 11.5 m above the ground.
a. calculate the following:
(i) the potential energy gained
(ii) the average power developed by the boy in walking up the stairs
b. As the boy slides down the slide and reaches the bottom, his speed is 12.0 m/s and the distance traveled along the slide is 108 m. Consider that there is friction present as the boy slides down.
(i) Find the mechanical energy lost by the boy in sliding down the slide.
(ii) What is the average frictional force acting on the boy?
1
Expert's answer
2018-11-05T15:52:09-0500

(i)

E=mgh=(50)(9.81)(11.5)=5640 J=5.46 kJ.

(ii)

P=E/t=5640/60=94.0 W.


b. As the boy slides down the slide and reaches the bottom, his speed is 12.0 m/s and the distance traveled along the slide is 108 m. Consider that there is friction present as the boy slides down.

(i) Find the mechanical energy lost by the boy in sliding down the slide.

(ii) What is the average frictional force acting on the boy?

Solution

(i)

E_lost=E-K=mgh-(mv^2)/2

E_lost=50((9.81)(11.5)-〖12〗^2/2)=2040 J=2.04 kJ

(ii)

F_fr=E_lost/v=2040/12=170 N.